OpenStudy (anonymous):

There exists a 3-dimensional structure such that at height h the cross section is a square with side length 7sinh+7. If the height of the structure is 4π, then the volume of the structure can be expressed as aπ, where a is a positive integer. What is the value of a?

OpenStudy (anonymous):

@waterineyes @AravindG @satellite73 @ganeshie8 @mathslover

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ash2326 @Mertsj @.Sam. @Callisto @amistre64

OpenStudy (ash2326):

Volume = Cross section Area * height here the cross section area is different at different height. Let at height h, we consider a small height dh the cross section area is \[(7\sin h+7)^2\] Do you understand this?

OpenStudy (anonymous):

why square?

OpenStudy (ash2326):

Square side is 7 sin h +7 so areas will be the square of the side

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

so the volume will be \[\int (7 \sin h +7)^2\times dh\] height varies from 0 to 4pi \[\int_{0}^{4\pi} (7 \sin h +7)^2\times dh\] Can you do this integration?

OpenStudy (anonymous):

what's the ans.?

OpenStudy (anonymous):

i got 294

OpenStudy (ash2326):

Yes, me too.

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

welcome :D